OpenStudy (anonymous):

Solve the equation for x. 3x =-4y

OpenStudy (anonymous):

Get x alone in one side by doing inverse operations to everything else in that side. In this problem for example, divide 3 from both sides.

OpenStudy (anonymous):

Move all terms containing x to the left, all other terms to the right. Divide each side by '3'. x = -1.333333333y

OpenStudy (anonymous):

i have no choices for those answers

OpenStudy (stamp):

3x = -4y Divide each side by three to obtain "x" from "3x" (3x)/3 = (-4y)/3 3/3 x = -4/3 y 1/1 x = -4/3 y x = -4/3 y

OpenStudy (anonymous):

thank u stamp

OpenStudy (anonymous):

Without simplifying the answer would be \[x=-4y/3\]

OpenStudy (anonymous):

x=-4y/3